A uniform 80.0 N ladder 4.0 m long is placed against a frictionless wall with its base situated 2.0 from the wall. Find the forces exerted by the wall and by the ground on the ladder.


1
Expert's answer
2021-02-18T18:48:02-0500

The equations of static equilibrium

become:

"F-N_2=0" (horizontal)


"N_1-mg=0(vertical)" and


"N_2Lsin\theta-\frac{1} {2} mgLcos\theta=0" (torque about the bottom of the ladder)


When "mg=80N, cos \theta=\frac{2.0}{4.0}=0.5"


"N_1" (the force exerted by the ground) =80N


"N_2" (the force exerted by the wall) ="\frac{1} {2} mgcos\theta=\frac{1} {2} ×80×0.5"


"N_2=20N"


F(The friction force exerted by the ground) ="N_2=20N."